A particle of mass m=1Kg moves in the x -y plane.The force on it at time t is
F(t)= [2sin(αt)i+3cos(αt)j]N
where α =1s^(-1) At time t=0 the particle is at rest at the origin. Calculate the magnitude of its position vector r (in m ) and velocity vector v(in m/s) at
time t=(π/2)s
Ans r=[(π-2)²+9]^(1/2)
V=√13
1
Expert's answer
2020-11-05T03:54:52-0500

Solution

Using newton equation

"m\frac{dV}{dt}="  [2sin(αt)i+3cos(αt)j]N

Integrate above equation with respect to T

"V= \frac{[-2cos(αt)i+3sin(αt)j]}{\alpha m}+c" .,,,,,,,,eq1


At initially t=0 V=0

We get c= -2i

So velcity become

"V= \frac{[-2cos(αt)i+3sin(αt)j]}{\alpha m}-2i"

Using "t=\pi/2" and "\alpha=1"

"V=-2i+3j"

Magnitude of this velocity is

"V=\sqrt{13} m/s"


Again integrate equation 1 we get

Position vector

"r= \frac{[-2sin(αt)i+3cos(αt)j]}{\alpha^2m}-2it+3j"

Putting all given value

Position vector become as


"r=[(π-2)²+9]^{\frac{1}{2}}m"